OpenStudy (anonymous):

Simplify. (x^-2x^-3)^4

OpenStudy (anonymous):

WOW! Haha that's a funny integral: \[\left(x^{-2x^{-3}}\right)^4=x^{-2\times4\times x^{-3}}=x^{-8x^{-3}}\]

OpenStudy (anonymous):